# Nickel and Diming Your GAN: A Dual-Method Approach to Enhancing GAN Efficiency via Knowledge Distillation

### v1

@@ 5 Limitations
Our method shows excellent performance via distribution matching, yet it tends to focus on the fidelity of generated images. In fact, every method experiences significant degradation in recall performance, even at low compression rates (Figure 4). Furthermore, there still remains the imbalance between the generator and discriminator at extreme compression rates, which incurs significant performance degradation. Thus, it is an interesting research direction to develop methods that are capable of maintaining diversity and stability when compressing generative models at extreme compression rates.

### v2

@@ 5 Limitations
Our method shows excellent performance via distribution matching, yet it tends to focus on the fidelity of generated images. In fact, every method experiences significant degradation in recall performance, even at low compression rates (Fig. 4). Furthermore, there still remains the imbalance between the generator and discriminator at extreme compression rates, which incurs significant performance degradation. Thus, it is an interesting research direction to develop methods that are capable of maintaining diversity and stability when compressing generative models at extreme compression rates.
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.2022R1C1C1008496), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2020-II201336, Artificial Intelligence graduate school support (UNIST), No.RS-2021-II212068, Artificial Intelligence Innovation Hub, RS-2022-II220959, (Part 2) Few-Shot Learning of Causal Inference in Vision and Language for Decision Making, RS-2022-II220264, Comprehensive Video Understanding and Generation with Knowledge-based Deep Logic Neural Network). We also thank the supercomputing resources of the UNIST Supercomputing Center.